# Energy exchange between electrons and ions in ion temperature gradient turbulence

### v1

@@ II.3 Spectral analysis of turbulent energy exchange in wavenumber space
We investigate the physical mechanism of the turbulent energy exchange by using Eq. (15). In the steady state of turbulence, the time derivative acting on the perturbed potential Eq. (5) can be transferred to that on the nonadiabatic perturbed distribution function with the help of the Leibniz rule. In addition, we use the gyrokinetic equation, Eq. (5), to rewrite Eq. (15) as
@@
$Q^{\rm turb}_{aC\bm{k}_{\perp}}$ $=$ $-\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}\int d^{3}vC_{a}\psi_{a\bm{k}_{\perp}}^{*}\Bigg{\rangle}\Bigg{\rangle}$ (28)
where $E_{r}=0$ is used as described after Eq. (II.2). The perturbed fields $\bm{G}_{a\bm{k}_{\perp}}$ and the perturbed currents $\bm{j}_{a\bm{k}_{\perp}}$ at the gyrocenter position are defined by
@@
Here, $\bm{j}_{a\psi\bm{k}_{\perp}}$ is derived from the nonlinear term of Eq. (5), and represents the current induced by the perturbed potential. As shown in Eq. (24), the turbulent energy exchange is caused by the product of the perturbed field and current due to the nonadiabatic distribution. The field and current can be decomposed into components in directions parallel and perpendicular to the background magnetic field. The perpendicular current component can also be classified as the two parts: the one is produced by the $\nabla B$-curvature drift in the toroidal magnetic field and the other by the drift due to the turbulent potential field. Furthermore, the effect of collisions is given by the last term at the right-hand side of Eq. (24), and thus the turbulent energy exchange is represented by the sum of the four parts.
@@
which can be illustrated by using the definition of $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ given in Eqs. (27), (29) and (34). Equation (36) implies that $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ does not contribute to the net heating or cooling of particles of the species $a$ but it represents the energy transfer in the wavenumber space through nonlinear interactions between different modes. Thus, $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ influences the profile of the total wavenumber spectrum $Q^{\rm turb}_{s\bm{k}_{\perp}}$. Because of Eq. (36), the total turbulent energy exchange $Q^{\rm turb}_{a}$ is given by taking the sum of the three components $Q^{\rm turb}_{a\parallel\bm{k}_{\perp}}$, $Q^{\rm turb}_{aB\bm{k}_{\perp}}$, and $Q^{\rm turb}_{aC\bm{k}_{\perp}}$ over the whole wavenumber space.
It is noted here that, using Eq. (5) and Eqs. (7)–(9), we can obtain
@@ II.4 Quasilinear model and entropy balance
In this section, the predictability of the quasilinear model for the turbulent energy exchange is discussed with the entropy balance in the linear and nonlinear states. When using the solutions $h_{a\bm{k}_{\perp}}$ and $\psi_{a\bm{k}_{\perp}}$ of the linearized version of Eq. (5) combined with Eqs. (7)–(9) to evaluate $Q^{\rm turb}_{a}$ defined in Eq. (15), we obtain
@@
The entropy balance equation for each wavenumber is derived from Eq. (5) as
@@
In the quasilinear model, the transport fluxes divided by the squared potential, $\Gamma^{\rm turb}_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ and $q^{\rm turb}_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, in the turbulent steady state are approximated by the corresponding values obtained from the linear analysis. We here include the turbulent energy exchange term into the quasilinear model and estimate $Y_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ from the linear calculation. When this model is valid, the values of the terms on the right-hand side of Eq. (46) divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ should not change whether linear or nonlinear simulations are performed to evaluate them. On the other hand, when divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, the time-derivative term and the nonlinear entropy transfer term on the left-hand side of Eq. (46), take different values between the linear and nonlinear cases. The time-derivative term is dominant and the nonlinear entropy transfer vanishes in the linear case, while the former is negligible and the latter is dominant in the nonlinear case. Here, the ratio of the collisional dissipation to the squared potential, $D_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, is assumed to take the same value. Then, it is concluded from the above-mentioned discussion of Eq. (46) based on the quasilinear model that the ratio of the time-derivative term divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ in the linear case should be the same as the value of $\mbox{}-N_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ in the nonlinear case in order to keep the balance between the left- and right-hand sides of Eq. (46).
@@
In Sec. III.5, the relative magnitude of each term in the entropy balance equation, Eq. (42), is evaluated by the linear and nonlinear gyrokinetic simulations and the speculations based on the quasilinear model described above are examined.
@@ III.1 Simulation settings
Figure 1: (a) The linear growth rate $\gamma$ and frequency $\omega_{r}$ at $(k_{x}\rho_{ti},k_{y}\rho_{ti})=(0,0.30)$, the turbulent ion heating $Q^{\rm turb}_{i}$, the turbulent particle flux $\Gamma^{\rm turb}_{e}=\Gamma^{\rm turb}_{i}$, and the turbulent heat fluxes $q^{\rm turb}_{s}$ $(s=e,i)$ plotted as functions of the temperature ratio $T_{e}/T_{i}$. (b) Comparison of all terms in the entropy balance equation, Eq. (II.2), in the saturated state of the ITG turbulence for $T_{e}/T_{i}=1.0$. All terms are normalized by the entropy production $q_{i}^{\rm turb}/{T_{i}L_{Ti}}$ due to the turbulent ion heat flux.
@@
In this research, microturbulence simulations are performed by the GKV codeGKV , which numerically solves the gyrokinetic equations in the flux-tube domain. We here focus on the Ion Temperature Gradient (ITG) mode turbulence so that the electron temperature gradient is set to zero, $R_{0}/L_{Te}=0$. Other plasma parameters used in the simulations are $R_{0}/L_{na}=2.22,R_{0}/L_{Ti}=6.92,m_{e}/m_{i}=5.446\times 10^{-4},e_{i}/e_{e}=-1,\epsilon_{r}=0.18,q_{0}=1.4,\hat{s}=0.78$, which are the same values as in the Cyclone D- III D base caseDimits . Here, $R_{0},\epsilon_{r},q_{0}$, and $\hat{s}=(r/q)(dq/dr)$ represent the major radius, inverse aspect ratio, safety factor, and magnetic shear, respectively. The flux tube coordinates $x=r-r_{0},y=r_{0}/q_{0}\left(q(r)\theta-\zeta\right),$ and $z=\theta$ are used in a low-$\beta$, large aspect ratio, axisymmetric torus with circular concentric flux surfaces where $r$, $\theta$, and $\zeta$ are minor radius, poloidal angle, and toroidal angle, respectively. The subscript 00 denotes parameters at the center of flux tube. The perpendicular wavenumber is given by $\bm{k}_{\perp}=\left(k_{x}+\hat{s}zk_{y}\right)\bm{e}_{r}+k_{y}\bm{e}_{\theta}$. Here, $k_{x}$ and $k_{y}$ are wavenumbers in the directions of $\nabla x$ and $\nabla y$, respectively, while $\bm{e}_{r}$ and $\bm{e}_{\theta}$ are unit vectors parallel to $\nabla r$ and $\nabla\theta$, respectively. The ion beta value is set to $\beta_{i}=4\pi n_{i}T_{i}/B^{2}=1\times 10^{-4}$ for which the electrostatic approximation is valid. In simulations performed here, $B_{\parallel\bm{k}_{\perp}}$ is neglected, although $A_{\parallel\bm{k}_{\perp}}$ is retained in order to avoid numerical difficulty due to very rapid electrostatic waves called the $\omega_{H}$ mode Lee . The Lenard-Bernstein collision operatorLenard is used here because it takes less computation time than more rigorous collision models. However, we expect that the collision model does not influence results from the present study where the normalized collision frequency is set to $\nu^{*}_{ii}\equiv R_{0}q\nu_{ii}/(\epsilon^{3/2}v_{ti})=0.068$, which is much smaller than the growth rates of the ITG modes in the present study. The temperature ratio is set in a range from $T_{e}/T_{i}=0.80$ to $1.5$.
For the nonlinear simulations, the domain sizes are basically set to $64.10\rho_{ti}\leq x\leq 64.10\rho_{ti},62.83\rho_{ti}\leq y\leq 62.83\rho_{ti},\pi\leq z\leq\pi,-4v_{ta}\leq v_{\parallel}\leq 4v_{ta}$, and $0\leq\mu B_{0}/T_{a}\equiv m_{a}v_{\perp}^{2}B_{0}/2T_{a}\leq 8$. Numbers of grid points in $(x,y,z,v_{\parallel},\mu)$ are $288\times 96\times 64\times 64\times 32$ or $288\times 96\times 256\times 64\times 32$. The resolved perpendicular wavenumbers are $-4.70\rho_{ti}^{-1}\leq k_{x}\leq 4.70\rho_{ti}^{-1}$ and $-1.55\rho_{ti}^{-1}\leq k_{y}\leq 1.55\rho_{ti}^{-1}$ with the minimum wavenumbers $(k_{x,\mathrm{min}},k_{y,\mathrm{min}})=(0.049\rho_{ti}^{-1},0.050\rho_{ti}^{-1})$. The output data shown in this paper are normalized by following Ref. GKV except for Fig. 2.
@@
Figure 2: Comparison of energy exchanges due to Coulomb collisions (blue squares) and turbulence (orange circles). The left figure shows the case of $\delta\equiv\rho_{ti}/R_{0}=9.6\times 10^{-4}$, $n_{e}=2.0\times 10^{19}[\rm{m^{-3}}]$, $T_{i}=0.9[\rm{keV}]$, and the right one shows the case of $\delta=1.7\times 10^{-3}$, $n_{e}=2.25\times 10^{20}[\rm{m^{-3}}]$, $T_{i}=3.0[\rm{keV}]$. Red triangles and black stars indicate turbulent energy exchanges obtained from simulations for $T_{e}/T_{i}=1.2$ and $R_{0}/L_{na}=3.0$ using different temperature gradients given by $(R_{0}/L_{Ti},R_{0}/L_{Te})=(5.0,0)$ and $(R_{0}/L_{Ti},R_{0}/L_{Te})=(5.0,6.5)$, respectively.
@@ III.2 Heat flux, turbulent energy exchange, and entropy balance
The ratio of each entropy balance term to the entropy production term caused by the ion heat flux $q^{\rm turb}_{i}/T_{i}L_{Ti}$ is shown in Fig. 1(b). The numerical error in the entropy balance defined by the difference between the left- and right-hand sides of Eq. (II.2) is within 6% of $q^{\rm turb}_{i}/T_{i}L_{Ti}$ for ions and 3% for electrons. In the case of ions, the particle and heat fluxes generate entropy while the turbulent energy exchange and collisions reduce entropy, thus maintaining balance. For electrons, on the other hand, the electron heat flux does not appear as an entropy-producing term because of no electron temperature gradient, although the turbulent energy exchange plays a major role in the entropy production in addition to the electron particle flux. The generated entropy is dissipated by collisions, and the entropy of turbulent fluctuations in the electron distribution function is kept in a steady state. The same result as described above is also reported in ref. Maeyama .

### v2

@@ II.3 Spectral analysis of turbulent energy exchange in wavenumber space
We investigate the physical mechanism of the turbulent energy exchange by using Eq. (15). In the steady state of turbulence, the time derivative acting on the perturbed potential Eq. (II.1) can be transferred to that on the nonadiabatic perturbed distribution function with the help of the Leibniz rule. As shown in Appendix B, using the gyrokinetic equation, Eq. (II.1), we can rewrite Eq. (15) as
@@
$Q^{\rm turb}_{aC\bm{k}_{\perp}}$ $=$ $-\operatorname{\rm{Re}}\Bigg{\langle}\Bigg{\langle}\int d^{3}vC_{a}^{GK}\psi_{a\bm{k}_{\perp}}^{*}\Bigg{\rangle}\Bigg{\rangle}$ (28)
where $E_{r}=0$ is used. The perturbed fields $\bm{G}_{a\bm{k}_{\perp}}$ and the perturbed currents $\bm{j}_{a\bm{k}_{\perp}}$ at the gyrocenter position are defined by
@@
Here, $\bm{j}_{a\psi\bm{k}_{\perp}}$ is derived from the nonlinear term of Eq. (II.1), and represents the current induced by the perturbed potential. As shown in Eq. (24), the turbulent energy exchange is caused by the product of the perturbed field and current due to the nonadiabatic distribution. The field and current can be decomposed into components in directions parallel and perpendicular to the background magnetic field. The perpendicular current component can also be classified as the two parts: the one is produced by the $\nabla B$-curvature drift in the toroidal magnetic field and the other by the drift due to the turbulent potential field.
Furthermore, the effect of collisions is given by the last term at the right-hand side of Eq. (24), and thus the turbulent energy exchange is represented by the sum of the four parts.
@@
which can be illustrated by using the definition of $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ given in Eqs. (27), (29) and (34). Equation (36) implies that $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ does not contribute to the net heating or cooling of particles of the species $a$ but it represents the energy transfer in the wavenumber space through nonlinear interactions between different modes. Thus, $Q^{\rm turb}_{a\psi\bm{k}_{\perp}}$ influences the profile of the total wavenumber spectrum $Q^{\rm turb}_{a\bm{k}_{\perp}}$. Because of Eq. (36), the total turbulent energy exchange $Q^{\rm turb}_{a}$ is given by taking the sum of the three components $Q^{\rm turb}_{a\parallel\bm{k}_{\perp}}$, $Q^{\rm turb}_{aB\bm{k}_{\perp}}$, and $Q^{\rm turb}_{aC\bm{k}_{\perp}}$ over the whole wavenumber space. In particular, $Q^{\rm turb}_{a\parallel\bm{k}_{\perp}}$, $Q^{\rm turb}_{aB\bm{k}_{\perp}}$ are the contributions to Joule heating (cooling) via currents parallel and perpendicular to the background magnetic field.
It is noted here that, using Eq. (II.1) and Eqs. (7)–(9), we can obtain
@@ II.4 Quasilinear model and entropy balance
In this section, the predictability of the quasilinear model for the turbulent energy exchange is discussed with the entropy balance in the linear and nonlinear states. When using the solutions $h_{a\bm{k}_{\perp}}$ and $\psi_{a\bm{k}_{\perp}}$ of the linearized version of Eq. (II.1) combined with Eqs. (7)–(9) to evaluate $Q^{\rm turb}_{a}$ defined in Eq. (15), we obtain
@@
The entropy balance equation for each wavenumber is derived from Eq. (II.1) as
@@
In the quasilinear model, the transport fluxes divided by the squared potential, $\Gamma^{\rm turb}_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ and $q^{\rm turb}_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, in the turbulent steady state are approximated by the corresponding values obtained from the linear analysis. We here include the turbulent energy exchange term into the quasilinear model and estimate $Y_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ from the linear calculation.
Table 1: Plasma and field parameters
Plasma and field parameters Value
Normalized ion temperature gradient $R_{0}/L_{Ti}$ 6.92
Normalized electron temperature gradient $R_{0}/L_{Te}$ 0
Normalized density gradient $R_{0}/L_{na}$ 2.22
Mass ratio $m_{e}/m_{i}$ $5.446\times 10^{-4}$
Charge ratio $e_{e}/e_{i}$1
Ion beta value $\beta_{i}=4\pi n_{i}T_{i}/B^{2}$ $1\times 10^{-4}$
Normalized collision frequency $\nu^{*}_{ii}\equiv R_{0}q_{0}\nu_{ii}/(\epsilon_{r}^{3/2}v_{ti})$ 0.068
Temperature ratio $T_{e}/T_{i}$ $0.8\sim 1.5$
Inverse aspect ratio $\varepsilon_{r}$ $0.18$
Safety factor $q_{0}$ $1.4$
Magnetic shear $\hat{s}=(r/q)(dq/dr)$ $0.78$
Table 2: Resolution settings
Domain sizes and
resolved perpendicular wavenumbers
$-64.10\rho_{ti}\leq x\leq 64.10\rho_{ti}$
$-62.83\rho_{ti}\leq y\leq 62.83\rho_{ti}$
$-\pi\leq z\leq\pi$
$-4v_{ta}\leq v_{\parallel}\leq 4v_{ta}$
$0\leq\mu B_{0}/T_{a}\equiv m_{a}v_{\perp}^{2}/2T_{a}\leq 8$
$-4.70\rho_{ti}^{-1}\leq k_{x}\leq 4.70\rho_{ti}^{-1},(k_{x,\mathrm{min}}=0.049\rho_{ti}^{-1})$
$-1.55\rho_{ti}^{-1}\leq k_{y}\leq 1.55\rho_{ti}^{-1},(k_{y,\mathrm{min}}=0.050\rho_{ti}^{-1})$
Grid points in $(x,y,z,v_{\parallel},\mu)$
$288\times 96\times 64\times 64\times 32$
$288\times 96\times 256\times 64\times 32$
When this model is valid, the values of the terms on the right-hand side of Eq. (46) divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ should not change whether linear or nonlinear simulations are performed to evaluate them. On the other hand, when divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, the time-derivative term and the nonlinear entropy transfer term on the left-hand side of Eq. (46), take different values between the linear and nonlinear cases. The time-derivative term is dominant and the nonlinear entropy transfer vanishes in the linear case, while the former is negligible and the latter is dominant in the nonlinear case. Here, the ratio of the collisional dissipation to the squared potential, $D_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$, is assumed to take the same value. Then, it is concluded from the above-mentioned discussion of Eq. (46) based on the quasilinear model that the ratio of the time-derivative term divided by $\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ in the linear case should be the same as the value of $\mbox{}-N_{a\bm{k}_{\perp}}/\langle\langle|\phi_{\bm{k}_{\perp}}|^{2}\rangle\rangle$ in the nonlinear case in order to keep the balance between the left- and right-hand sides of Eq. (46).
@@
The discussion of quasilinear weights with entropy balance is expected to apply not only to tokamaks but also to more generic 3D geometries. In Sec. III.5, the relative magnitude of each term in the entropy balance equation, Eq. (42), is evaluated by the linear and nonlinear gyrokinetic simulations, and the speculations about the quasilinear model described above are examined. It should be pointed out that in this work, we discuss only quasilinear weights and do not investigate the saturation rule required to predict fluxes.
@@ III.1 Simulation settings
Figure 1: (a) The linear growth rate $\gamma$ and frequency $\omega_{r}$ at $(k_{x}\rho_{ti},k_{y}\rho_{ti})=(0,0.30)$, the turbulent ion heating $Q^{\rm turb}_{i}$, the turbulent particle flux $\Gamma^{\rm turb}_{e}=\Gamma^{\rm turb}_{i}$, and the turbulent heat fluxes $q^{\rm turb}_{a}$ $(a=e,i)$ plotted as functions of the temperature ratio $T_{e}/T_{i}$. (b) Comparison of all terms in the entropy balance equation, Eq. (12), in the saturated state of the ITG turbulence for $T_{e}/T_{i}=1.0$. All terms are normalized by the entropy production $q_{i}^{\rm turb}/{T_{i}L_{Ti}}$ due to the turbulent ion heat flux.
In this research, microturbulence simulations are performed by the GKV codeGKV , which solves the gyrokientic equation for the perturbed distribution function based on the Eulerian scheme in $(k_{x},k_{y},z,v_{\parallel},\mu)$ space. The nonlinear term is evaluated in the real space, and is transformed back to the wavenumber space by means of 2D Fast Fourier Transform algorithm and the 2/3 rule in $(k_{x},k_{y})$. It employs the local flux-tube domain where the background densities, temperatures, and their gradients are fixed. While we do not deal with 3D geometries here, microturbulence in herical systems has also been studied by GKVherical ; herical1 .
The flux tube coordinates for a low-$\beta$, large aspect ratio, axisymmetric torus with concentric circular cross sections, $x=r-r_{0},y=r_{0}/q_{0}\left(q(r)\theta-\zeta\right),$ and $z=\theta$ are used in this work, where $r$, $\theta$, and $\zeta$ are minor radius, poloidal angle, and toroidal angle, respectively. The subscript 00 denotes parameters at the center of flux tube. The perpendicular wavenumber is given by $\bm{k}_{\perp}=\left(k_{x}+\hat{s}zk_{y}\right)\bm{e}_{r}+k_{y}\bm{e}_{\theta}$. Here, $k_{x}$ and $k_{y}$ are wavenumbers in the directions of $\nabla x$ and $\nabla y$, respectively, while $\bm{e}_{r}$ and $\bm{e}_{\theta}$ are unit vectors parallel to $\nabla r$ and $\nabla\theta$, respectively.
@@
We here focus on the Ion Temperature Gradient (ITG) mode turbulence in tokamak plasmas so that the electron temperature gradient is set to zero, $R_{0}/L_{Te}=0$, where $R_{0}$ represent the major radius. Plasma and field parameters used in the simulations are shown in Tab. LABEL:tab:Plasma_and_field_parameters. Most of them are the same values as in the Cyclone DIII-D base caseDimits . The ion beta value is set to $\beta_{i}=1\times 10^{-4}$ for which the electrostatic approximation is valid. In simulations performed here, $B_{\parallel\bm{k}_{\perp}}$ is neglected, although $A_{\parallel\bm{k}_{\perp}}$ is retained in order to avoid numerical difficulty due to very rapid electrostatic waves called the $\omega_{H}$ mode Lee . The Lenard-Bernstein collision operatorLenard is used here because it takes less computation time than more rigorous collision models. However, we expect that the collision model does not influence results from the present study where the normalized collision frequency is set to $\nu^{*}_{ii}\equiv R_{0}q\nu_{ii}/(\epsilon^{3/2}v_{ti})=0.068$, which is much smaller than the growth rates of the ITG modes in the present study. Since collisional energy exchange is proportional to the temperature difference between electrons and ions, the temperature ratio is set in a range from $T_{e}/T_{i}=0.80$ to $1.5$ in order to compare turbulent and collisional energy exchanges.
The resolution settings are shown in Tab. LABEL:tab:resolution. High resolution for the parallel coordinate $z$ is used in Fig. 1 (b) and Fig. 4 to suppress an error of entropy balances by a hyper diffusion termHyper Diffusion . There is no significant difference in turbulent fluxes and energy exchange when using either low or high resolutions for $z$. The output data shown in this paper are normalized by following Ref. GKV except for Fig. 2.
@@
Figure 2: Comparison of energy exchanges due to Coulomb collisions (blue squares) and turbulence (orange circles). The left figure (a) shows the case of $\delta\equiv\rho_{ti}/R_{0}=9.6\times 10^{-4}$, $n_{e}=2.0\times 10^{19}[\rm{m^{-3}}]$, $T_{i}=0.9[\rm{keV}]$, and the right one (b) shows the case of $\delta=1.7\times 10^{-3}$, $n_{e}=2.25\times 10^{20}[\rm{m^{-3}}]$, $T_{i}=3.0[\rm{keV}]$. Red triangles and black stars indicate turbulent energy exchanges obtained from simulations for $T_{e}/T_{i}=1.2$ and $R_{0}/L_{na}=3.0$ using different temperature gradients given by $(R_{0}/L_{Ti},R_{0}/L_{Te})=(5.0,0)$ and $(R_{0}/L_{Ti},R_{0}/L_{Te})=(5.0,6.5)$, respectively.
@@ III.2 Heat flux, turbulent energy exchange, and entropy balance
The ratio of each entropy balance term to the entropy production term caused by the ion heat flux $q^{\rm turb}_{i}/T_{i}L_{Ti}$ is shown in Fig. 1(b). The numerical error in the entropy balance defined by the difference between the left- and right-hand sides of Eq. (12) is within 6% of $q^{\rm turb}_{i}/T_{i}L_{Ti}$ for ions and 3% for electrons. In the case of ions, the particle and heat fluxes generate entropy while the turbulent energy exchange and collisions reduce entropy, thus maintaining balance. For electrons, on the other hand, the electron heat flux does not appear as an entropy-producing term because of no electron temperature gradient, although the turbulent energy exchange plays a major role in the entropy production in addition to the electron particle flux. The generated entropy is dissipated by collisions, and the entropy of turbulent fluctuations in the electron distribution function is kept in a steady state. The same result as described above is also reported in ref. Maeyama .